OpenStudy (epicteatime):

What is the mass, in grams, of a sample of 1.72 × 10^23 atoms of potassium (K)? Show your work or explain the steps that you used to determine your answer. i don't know how i keep getting 2.86...mol K!!

OpenStudy (anonymous):

First, you need to convert atoms to moles; so \[1.72\times10^{23}\times \left( \frac{ 1 mol }{ 6.022\times10^{23}} \right)\], which equals: 0.2856 mols. Then you convert to grams by multiplying by the atomic mass number.\[0.2856\times \left( \frac{ 39g K }{ 1 mol K} \right)\] which equals 11.1

OpenStudy (epicteatime):

thanks mate!!